OpenStudy (anonymous):

An investment is advertised as returning 4.5% every 6 months (semiannually), compounded semiannually. If $30,000 is invested, the growth can be modeled by the equation A(t) = 30,000(1.045)2t. What is the equivalent annual growth rate for this investment (rounded to the nearest tenth of a percent) and what is it worth (rounded to the nearest thousand dollar) after 30 years? Hint: Find the value of 1.0452 on your calculator.

OpenStudy (mathmale):

It's really important that you learn to express exponentiation properly. Your A(t) = 30,000(1.045)2t seems to imply multiplication all the way through, which is not correct. Please use Equation Editor or the Draw utility, or the ' ^ ' symbol, to express exponentiation. \[Right: A(t) = 30,000(1.045)^{2t}\] Wrong: A(t) = 30,000(1.045)2t

OpenStudy (mathmale):

If you wish to take the hint given you, evaluate (1.045)^2 on your calculator. Type in 1.045. Then press the calculator key that looks like ^ . Your answer?

OpenStudy (anonymous):

Hi there, yes I see what you mean

OpenStudy (anonymous):

i have to show the exponents correctly, thank you for the tip!